Quarterly billing for water usage is shown below.

Year

Quarter

1

2

3

4

Winter

64

66

68

73

Spring

103

103

104

120

Summer

152

160

162

176

Fall

73

72

78

88

The forecast equation is: Ft = 64.875 – 6.1375 Qtr1t + 32.325 Qtr2t + 86.0375 Qtr3t + 1.2875 (t). Forecast the summer of year 5 and spring of year 6.
Answer #1

from given forecast equation:

Forecast for summer of year 5 =64.875 +86.0375+1.2875*19 =175.3750

Forecast for spring of year 6 =64.875 +32.325+1.2875*22 =125.525